For a parabolic dish antenna, show that the
–3db beam width with D = 6m and Wavelength= 3 cm is 0.35°.
1
Expert's answer
2019-06-10T09:41:56-0400

The beamwidth of the parabolic dish antenna can be estimated as


"\Psi=\frac{70\lambda}{D}=\frac{70\cdot0.03}{6}=0.35^\circ."

At


"10\text{log}\frac{1}{2}=-3\text{ db}"

the power becomes twice as less.